The length of a rectangular window is 5 feet more than the with. The area of the window is 36 square feet. Find the length and with of window
Answer by fractalier(6550)

By inspection I can see it must be a 4' x 9' window, but we solve it by
L = W + 5
and
A = LW = 36
(W+5)(W) = 36
W^2 + 5W - 36 = 0
(W + 9)(W - 4) = 0
W = 4 feet and then
L = W + 5 = 9 feet